OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and has a slope of –1 in point-slope form. x – 4 = –1(y – 2) x – 2 = –1(y – 4) y – 2 = –1(x – 4) y – 4 = –1(x – 2)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

Wait, which one would it be? The last three have the same slope and intersect (2,4)

OpenStudy (anonymous):

According to my graphs atleast

pooja195 (pooja195):

Well i gave u the equation u should be able to do it from here

OpenStudy (anonymous):

y - 4 = -1(x -2)! ?

OpenStudy (anonymous):

If thats right how do I medal you?

OpenStudy (anonymous):

Yo

pooja195 (pooja195):

Yes thats right :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

Thanks man! I was confused on it for a while and scramble through my book for it.

pooja195 (pooja195):

im a girl ;)

OpenStudy (anonymous):

Oh my bad, Gal!